OpenStudy (anonymous):

Help! WILL FAN AND MEDAL! In a candy box, there are 4 times as many candies with nuts as candies without nuts. There are 25 candies in the box. How many candies with nuts are in the box? 20 9 5 21

OpenStudy (anonymous):

4x+x=25

OpenStudy (anonymous):

4(20)+20=100, so no 4(9)+9=45, so no again 4(5)+5=25, so yes!

OpenStudy (anonymous):

5?

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

25 is the answer?

OpenStudy (anonymous):

no, 5 is the answer

OpenStudy (anonymous):

oh lol! Thank-You

OpenStudy (anonymous):

Sorry, but 20 was the answer.